how do i find the squares?
(7p+6q)^2

Answer by uma(370)

(7p+6q)^2
Method 1: Using formula
(a+b)^2 = a^2+2ab+b^2
(7p+6q)^2 = (7p)^2+2(7p)(6q)+(6q)^2
= 49p^2 + 84pq + 36q^2

Method 2: By multiplying
(7p+6q)^2 = (7p+6q)(7p+6q)
= 7p(7p) + 7p(6q) + 6q(7p) +(6q)(6q)
= 49p^2 + 42pq + 42pq + 36q^2
= 49p^2 + 84pq + 36q^2


Good Luck!!!